OpenStudy (anonymous):

Solve the integral: 4x / (x^2 + 1) Show steps.

OpenStudy (anonymous):

\[u=x^{2}+1 .......du=2xdx......\]

OpenStudy (anonymous):

\[\int\limits_{}^{}2\frac{du}{u}\]

OpenStudy (anonymous):

=2 lnu +C substitute u back in and its finished

OpenStudy (anonymous):

\[=2\ln(x^{2}+1) +C\]